# PAX-TS: Model-agnostic multi-granular explanations for time series forecasting via localized perturbations

### v1

@@ III-C Trend-drift adjustment
$\mathbf{x}^{\prime}\;=\;\mathbf{x}_{i}-a+a^{\prime}-(m-m^{\prime})(i-1),\qquad\forall\,i\in\{1,\dots,b\}$
III-D Structured perturbation analysis
@@
Algorithm 1 describes how to apply PAX-TS for perturbation analysis. Given a set of scale parameters $\mathcal{A}$, and following the equations of the preceding sections, various characteristics of a time series subsequence $\mathbf{x}$ can be perturbed. In combination with a trained forecasting model $M_{\theta}(\cdot)$, forecasts $\hat{\mathbf{y}}^{\prime}$ can be made for the perturbed input $\mathbf{x}^{\prime}$. Next, a property of interest function $\pi\in\mathcal{P}$ is used, which extracts a human-comprehensible property (such as trend, maximum, or the value at a given index) of a time series. These functions are typically simple, such as $\max(x)$ or retrieving the value $x_{t}$, and we do not describe them in further detail. Given the distance $\Delta$ between the perturbed input and the original input, the change ratio $r_{\pi\alpha}$ between a property of interest of the original prediction $\hat{\mathbf{y}}$ and the perturbed prediction $\hat{\mathbf{y}}^{\prime}$ can be calculated. Using this change ratio, PAX-TS can quantify the effect of applying the perturbation $p(\cdot)$ on the input with relation to a given property of interest $\pi(\cdot)$ of the forecast. When averaging over all scale parameters in $\mathcal{A}$, the mean change ratio $\bar{r}_{\pi}$ can be calculated, while taking the sign of $\alpha$ into account, as negative scale parameters produce inverse change ratios. Algorithm 1 can be applied both to an individual forecasting sample and to an entire dataset by substituting $\mathbf{x}$.
Applying PAX-TS for multiple properties of interest on both the input-level and the forecast-level results in a matrix of change ratios, which can be visualized to illustrate and explain the behavior of the forecasting model $M_{\theta}(\cdot)$, as seen in Fig. 2. With this type of explanation, PAX-TS can answer end-user questions on summary statistics, extrema, and trend of the input and forecast. For example, the figure shows a negative correlation between the input minimum and the output’s maximum, mean, and variance. If a user aims to increase the forecast maximum, adjusting the input’s mean, variance, and trend appears to be more effective than increasing its maximum value.
III-E Complexity Analysis
To calculate change ratios with PAX-TS as outlined in Algorithm 1, $|\mathcal{A}|+1$ inference operations need to be performed, which is the most computationally complex part of the method. Additional necessary operations are perturbation, distance calculation, and property of interest extraction. However, as these operations all scale with $O(b)$ and are, therefore, less computationally intensive than the inference operations, they do not contribute to the overall complexity of applying our method. This shows that our method scales linearly with the number of scale parameters $|\mathcal{A}|$, which can be increased to improve the robustness of the method. In practice, $|\mathcal{A}|$ can be set to a low number, e.g. $\leq 4$. As suggested in early experiments, change ratios tend to be constant for different scale parameters when normalized by the distance, so that
$r_{\pi\alpha_{1}}+r_{\pi\alpha_{2}}\leq\epsilon.$
III-F Multivariate cross-channel correlation analysis
@@
PAX-TS enables the analysis of multivariate time series data to assess how a forecasting model takes cross-channel correlations into consideration. This can be done by slightly adapting the procedure of Algorithm 1 to highlight channel correlations of multivariate data. Algorithm 2 illustrates this in detail. Each time series channel is locally perturbed at each input time step (line 7). This allows the computation of the change ratio $r_{\pi ct\alpha}$, with respect to the output time step $\pi$, channel $c$, input time step $t$, and scale parameter $\alpha$. When averaging across all input and output time steps, we can calculate $\bar{R}_{c}$, which is the average change ratio of one channel with respect to all other channels of the time series. Finally, when aggregating the results for all channels, this results in $\bar{R}$, a $d\times d$ matrix of cross-channel correlations. We chose to calculate cross-channel correlations by perturbing all time steps of the input, so that $t\in\{1,\dots,b\}$, to more accurately capture the forecaster’s behavior and account for potential seasonal effects. For datasets with a large number of channels, the computational complexity of multivariate cross-channel correlation can be reduced by setting the number of scale parameters to a low value (e.g., $\alpha=2$), and timesteps in the input can be changed in a sparsified manner (e.g. $t\in\{1,3,5,\dots,n\}$. However, in the datasets we analyzed in this paper, a reduction was not necessary, as Algorithm 2 generally finished in less than 1 minute per dataset.
IV Experiments
Algorithm Ref. Approach
Naïve - Naïve
DLinear [22] MLP
MultiPatchFormer [3] Transformer
SegRNN [7] RNN
TimeMixer [8] MLP
iTransformer [2] Transformer
TimesFM [4] LLM
TABLE I: Algorithms evaluated in this study.
Dataset Ref. Domain Frequency Variates
M4 Hourly [23] Mixed Hourly 1
Weather (a){(a)} Weather Daily 1
Transactions [24] Sales Daily 1
CIF [25] Finance Monthly 1
Rain (b){(b)} Weather Monthly 1
M4 Yearly [23] Mixed Yearly 1
Covid (c){(c)} Healthcare Weekly 6
ETTh1 [26] Electricity Hourly 7
Exchange Rate [27] Finance Daily 7
Illness (d){(d)} Healthcare Weekly 7
TABLE II: Datasets used in this study. (a) https://www.bgc-jena.mpg.de/wetter/, (b) https://www.kaggle.com/datasets/macaronimutton/mumbai-rainfall-data, (c) https://www.kaggle.com/datasets/mexwell/sars-cov-2-germany, (d) https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
IV-A Experimental Setup
We test our proposed method with seven state-of-the-art forecasting algorithms (Table I) on ten different datasets (Table II), both to demonstrate that the method is model-agnostic and to investigate and explain different types of algorithms for different types of forecasting scenarios. All experiments were conducted on a single Nvidia A100 GPU, and results were averaged over three random states. We set the forecasting horizon length to $h=20$ and input length to $b=20$ in order to demonstrate a short-term forecasting scenario, where explainability can be highlighted more easily. However, our approach is applicable irrespective of the forecasting scenario and the values of these parameters. We use a temporally ordered, non-overlapping train-validation-test split of 7-1-2, and scale parameters $\mathcal{A}=\{-0.1,-0.05,-0.01,0.01,0.05,0.1\}$ to provide more robust results. We set $w=2$, $s=1$, configure $k$ according to the dataset, and we use left-fixed trend adjustment, which provided the best results during early experiments. We scale all inputs to have zero mean and unit variance and consider three error metrics: The Mean Absolute Error (MAE), and the Mean Squared Error (MSE), which are standard in multivariate forecasting evaluation as well as the Overall Weighted Average (OWA) [23], which is a combination of Symmetric Mean Absolute Percentage Error (sMAPE) and Mean Absolute Scaled Error (MASE), commonly employed for short-term forecasting evaluation. All experiments and results are publicly available on our anonymous git repository [^2] https://anonymous.4open.science/r/pax-ts-6410, where we share additional insights on PAX-TS and its explanations.
IV-B Forecasting performance
We conducted a benchmark on the forecasting performance of all models on all datasets, which is presented in Table IV on the last page to facilitate readability. The table additionally shows temporal dependency patterns, which we will present in Sec. IV-D. Notably, TimeMixer, MultiPatchFormer, and iTransformer perform best, with a significant Bonferroni-Dunn test [28] compared to the Naïve model. On the multivariate datasets Covid, Exchange Rate, and Illness, none of the benchmarked models achieves considerably better performance than the naïve forecaster, highlighting the difficulty of the multivariate forecasting problem. We additionally measured the total inference time across all datasets (Algorithm 1 line (2) + line (6)), where the LLM TimesFM incurred a significantly higher time (1390s) compared to TimeMixer (20.57s), MultiPatchFormer (8.89s), and iTransformer (1.08s).
IV-C High granularity: Time step importance
(a) SegRNN: Last-time step
(b) MultiPatchFormer: Diagonals
Pattern Properties # of occurrences Error $e_{norm}$
Other Other patterns, typically random and noisy. 5 0.7149
TABLE III: Overview of temporal dependency patterns and their associated error, averaged for all algorithms across all datasets.
PAX-TS can be applied on a time-step-level at the highest granularity to highlight the importance of every time step of the input window with relation to a given property of interest of the forecast. For example, Fig. 3 shows an explanation of MultiPatchFormer’s forecast on a sample of the CIF dataset. Specifically, PAX-TS calculates the relative importance $\bar{r}_{\pi}$ of each input time step for the maximum of the forecast, where $\pi(\cdot)=Max$. In the shown example, the first 19 time steps of the input have a positive correlation ($\bar{r}_{\pi}>0$) with the maximum, while the last time step has a negative correlation ($\bar{r}_{\pi}<0$). With this type of explanation, the end-user can understand which time steps of the data are important, and how the model’s predicted maximum changes for different inputs. Setting $\pi(\cdot)=Max$ is an example - when using PAX-TS in practice, other properties of interest can be analyzed instead.
IV-D Medium granularity: Temporal dependencies
To analyze temporal dependencies at a medium level of granularity, and characterize the importance of all time steps of the input in relation to all time steps of the forecast, we set $\mathcal{P}=\{\hat{\mathbf{y}}^{\prime}_{1},\dots,\hat{\mathbf{y}}^{\prime}_{h}\}$ and apply PAX-TS for all time steps of the input, where $p(\cdot)$ is applied to $\mathbf{x}$ at each index $i\in\{1,\dots,b\}$. The resulting matrix of relative importance can be visualized to explain temporal dependencies, as seen in Fig. 4. The figures visualize the correlation of time steps for SegRNN and MultiPatchFormer on the Rain dataset. On this dataset, MultiPatchFormer reaches the highest performance, while SegRNN performs considerably worse. The time step correlation heatmaps underline this: SegRNN bases its forecast mostly on the last time step, giving little importance to previous time steps of the input. MultiPatchFormer, on the other hand, shows a clear seasonal correlation between time steps, visualized as diagonals spanning the entire matrix. As Rain consists of monthly data, the dataset has a seasonality period of $k=12$, which is, for example, explained by the high positive correlation between input time steps 16 and 4, with forecast time steps 20 and 8. These clearly visible patterns illustrate the model’s behavior to make a forecast, and are characteristic of specific model architectures, datasets, and forecasting performance. We investigated this in more detail for all benchmarked algorithms and datasets, and derived six classes of temporal dependency patterns, which are shown in Table III. Table IV additionally shows which algorithm-dataset pair resulted in which temporal dependency pattern. Additionally, we calculated the average error across all datasets as a normalized mean of MAE and MSE compared to the naïve forecast as follows:
IV-E Low granularity: Multivariate cross-channel correlations
At the lowest granularity, PAX-TS, can assess and visualize cross-channel correlations in multivariate time series forecasting scenarios. This can be done by only perturbing a specific channel of the input and calculating the mean change ratio for all channels of the output separately, as presented in Algorithm 2. Fig. 5 shows the cross-channel correlations, as observed for iTransformer on the ETTh1 dataset, which consists of seven channels. We selected this example, as ETTh1 is the multivariate dataset with the lowest OWA error, where all models perform relatively well. In the visualized explanation, stronger correlations are highlighted with higher line width and more intense color. The model considers two strong cross-channel correlations: Channel 5 affects both channels 0 and 2. These correlations are even stronger than the autocorrelation of channels 0 and 2. Further, we found that on the benchmarked datasets, iTransformer and MultiPatchFormer most frequently take cross-channel correlations into account.

### v2

@@ III-C Trend-drift adjustment
$\mathbf{x}^{\prime}\;=\;\{\,x_{i}^{\prime}\,\}_{i=1}^{b},\qquad x_{i}^{\prime}\;=\;x_{i}\;-\;(m-m^{\prime})(i-1),\quad\forall i\in\{1,\dots,b\}.$
III-D Structured perturbation analysis with PAX-TS
[…]
III-E Complexity Analysis
To calculate change ratios with PAX-TS, $|\mathcal{A}|+1$ inference operations need to be performed, which is the most computationally complex part of the method. Additional necessary operations are perturbation, distance calculation, and property of interest extraction. However, as these operations all scale with $O(b)$ and are, therefore, less computationally intensive than the inference operations, they do not contribute to the overall complexity of applying our method. This shows that our method scales linearly with the number of scale parameters $|\mathcal{A}|$, which can be increased to improve the robustness of the method. In practice, $|\mathcal{A}|$ can be set to a low number, e.g. $\leq 4$. As suggested in early experiments, change ratios tend to be constant for different scale parameters when normalized by the distance, so that
$r_{\pi\alpha_{1}}+r_{\pi\alpha_{2}}\leq\epsilon.$
III-F Multivariate cross-channel correlation analysis
PAX-TS enables the analysis of multivariate time series data to assess how a forecasting model takes cross-channel correlations into consideration. To do this, each time series channel is locally perturbed at each input time step. This allows the computation of the change ratio $r_{\pi ct\alpha}$, with respect to the output time step $\pi$, channel $c$, input time step $t$, and scale parameter $\alpha$. When averaging across all input and output time steps, we can calculate $\bar{R}_{c}$, which is the average change ratio of one channel with respect to all other channels of the time series. Finally, when aggregating the results for all channels, this results in $\bar{R}$, a $d\times d$ matrix of cross-channel correlations. We chose to calculate cross-channel correlations by perturbing all time steps of the input, so that $t\in\{1,\dots,b\}$, to more accurately capture the forecaster’s behavior and account for potential seasonal effects. For datasets with a large number of channels, the computational complexity of multivariate cross-channel correlation can be reduced by setting the number of scale parameters to a low value (e.g., $\alpha=2$), and timesteps in the input can be changed in a sparsified manner (e.g. $t\in\{1,3,5,\dots,n\}$). However, in the datasets we analyzed in this paper, a reduction was not necessary, as the algorithm generally finished in less than 1 minute per dataset.
IV Experiments
IV-A Experimental Setup
We test our proposed method with seven forecasting algorithms (see Table III in the appendix) on ten datasets, including both univariate and multivariate datasets (see Table IV in the appendix), both to demonstrate that the method is model-agnostic and to investigate and explain different types of algorithms for different types of forecasting scenarios. We set the forecasting horizon length to $h=20$ and input length to $b=20$ in order to demonstrate a short-term forecasting scenario, where explainability can be highlighted more easily. However, our approach is applicable irrespective of the forecasting scenario and the values of these parameters. We use a temporally ordered, non-overlapping train-validation-test split of 7-1-2, and scale parameters $\mathcal{A}=\{-0.1,-0.05,-0.01,0.01,0.05,0.1\}$ to provide more robust results. We set $w=2$, $s=1$, and configure $k$ according to the dataset. We consider three error metrics: The Mean Absolute Error (MAE), and the Mean Squared Error (MSE), which are standard in multivariate forecasting evaluation as well as the Overall Weighted Average (OWA) [makridakis2020m4], which is commonly employed for short-term forecasting evaluation. All experiments and results are publicly available on our anonymous git repository [^2] https://anonymous.4open.science/r/pax-ts-6410, where we share additional insights on PAX-TS and its explanations. We conducted a benchmark on the forecasting performance of all models on all datasets, which is presented in Table VI in the appendix to facilitate readability. The table additionally shows temporal dependency patterns, which we will present in Sec. IV-C.
IV-B High granularity: Time step importance
PAX-TS can be applied on a time-step-level at the highest granularity to highlight the importance of every time step of the input window with relation to a given property of interest of the forecast. For example, Fig. 2 shows an explanation of MultiPatchFormer’s (see [naghashi2025multiscale]) forecast on a sample of the CIF dataset. Specifically, PAX-TS calculates the relative importance $\bar{r}_{\pi}$ of each input time step for the maximum of the forecast, where $\pi(\cdot)=max$. In the shown example, the first 19 time steps of the input have a positive correlation ($\bar{r}_{\pi}>0$) with the maximum, while the last time step has a negative correlation ($\bar{r}_{\pi}<0$). With this type of explanation, the end-user can understand which time steps of the data are important, and how the model’s predicted maximum changes for different inputs. Setting $\pi(\cdot)=Max$ is an example - when using PAX-TS in practice, other properties of interest can be analyzed instead.
IV-C Medium granularity: Temporal dependencies
(a) SegRNN: Last-time step
(b) MultiPatchFormer: Diagonals
Temporal dependencies at medium granularity are analyzed by using a set of properties of interest $\mathcal{P}=\{\pi_{i}\}_{i=1}^{h}$, where $\pi_{i}=\hat{y}^{\prime}_{i}$ and applying PAX-TS across all input indices $i\in\{1,\dots,b\}$, where $p(\cdot)$ perturbs $\mathbf{x}$ at each $i$. The resulting relative-importance matrix visualizes dependencies between input and forecast time steps (Fig. 3). On Rain, MultiPatchFormer reaches the best accuracy, whereas SegRNN ([lin2023segrnn]) performs considerably worse. The heatmaps reflect this: SegRNN relies almost exclusively on the last input step, assigning little weight to earlier steps, while MultiPatchFormer exhibits clear seasonal diagonals across the matrix. Because Rain is monthly data, the seasonality is $k=12$, shown by strong positive correlations between, for example, input index 16 and forecast index 20. These patterns characterize model behavior, dataset properties, and forecasting quality. Across all benchmarked algorithms and datasets, six temporal-dependency classes were derived (Table II), which are shown per-dataset in Table VI. Finally, the average error across datasets was computed as a normalized mean of MAE and MSE relative to the naïve forecast as follows:
Pattern Properties # of occurrences Error $e_{norm}$
Other Other patterns, typically random and noisy. 5 0.7149
TABLE II: Overview of temporal dependency patterns and their associated error, averaged for all algorithms across all datasets.
IV-D Low granularity: Multivariate cross-channel correlations
At the lowest granularity, PAX-TS, can assess and visualize cross-channel correlations in multivariate time series forecasting scenarios. This can be done by only perturbing a specific channel of the input and calculating the mean change ratio for all channels of the output separately. Fig. 4 shows the cross-channel correlations, as observed for iTransformer on the ETTh1 dataset, which consists of seven channels. We selected this example, as ETTh1 is the multivariate dataset with the lowest OWA error, where all models perform relatively well. In the visualized explanation, stronger correlations are highlighted with higher line width and more intense color. The model considers two strong cross-channel correlations: Channel 5 affects both channels 0 and 2. These correlations are even stronger than the autocorrelation of channels 0 and 2. Further, we found that on the benchmarked datasets, iTransformer and MultiPatchFormer most frequently utilize cross-channel correlations.
IV-E Comparison to ShapTime and TS-MULE
(a) ShapTime
(b) TS-MULE
We compare the explanations generated by PAX-TS with two explainability algorithms for forecasting: TS-MULE [schlegel2021ts] and ShapTime [zhang2023shaptime]. Fig. 5 shows explanations as generated by the two methods, following the visualization style of the original papers, for the same sample of the CIF dataset previously shown in Fig. 2, where we highlighted high-granularity time step importance in relation to the maximum of the forecast. Both ShapTime and TS-MULE are not able to capture importance for specific properties of interest, such as the maximum of the forecast, while PAX-TS can derive importance scores for different properties of interest, including extrema, summary statistics, and values at a given index. Further, TS-MULE’s results are always based on segment importance, rather than timestep importance, having a lower granularity than PAX-TS. The granularity of ShapTime can be configured with the number of segments, up to the time step level. However, as ShapTime evaluates all possible distinct subsets, this results in a time complexity of $O(2^{b})$ inference operations, compared to the $O(|\mathcal{A}|b)$ inference operations of PAX-TS to generate importance scores at the time-step-level. Further, given a trained model, both ShapTime and PAX-TS are deterministic, giving them high stability, while TS-MULE produces random results with high variance, leading to high instability. Since ShapTime observes changes in the mean of the forecast, the resulting importance scores are similar to PAX-TS, when setting $\pi(\cdot)=Mean$. Both ShapTime and TS-MULE are not suitable for application to multivariate data, whereas PAX-TS can be used to analyze cross-channel correlations, which produces a correlation graph, as shown in Fig. 4. All explanations produced by PAX-TS are directly based on input-output mappings of the forecasting algorithm, giving the explanations a high degree of fidelity, which we empirically prove in the appendix. Both TS-MULE and ShapTime operate on a local scope, giving explanations for an individual forecasting sample. PAX-TS, on the other hand, can produce both local and global explanations, increasing its scope of applicability and representativeness.
V Discussion
PAX-TS can generate explanations for time series forecasting models at different granularities, which are suitable for different use cases, depending on the aim of the end-user. In Sec. I, we introduced typical questions from end-users, which we relate to PAX-TS’ explanations here.
“Why is the forecast of productivity at 15:00 so low?”
“How can we increase the trend of the forecast?”
Lastly, PAX-TS can answer questions on cross-channel correlations of multivariate data with a graph, as shown with the example in Fig. 4. For more specific queries between channels, time-step-level index correlation can be used to illustrate the relation between the channels on a higher granularity. To apply PAX-TS, a top-down approach with increasingly higher explanation granularity can be employed: For multivariate data, first, a cross-channel correlation graph can be used to unveil correlations between different channels of the data. With this information, a heatmap of index correlations can be visualized to illustrate the correlation between two channels. When the end-user investigates a specific time step, extremum, or summary statistic, a stemplot can be generated to assess the time step importance of the specified sample for the given property of interest. With this highest level of granularity, PAX-TS can assist the end-user in comprehending time-step-level importance of the data.
PAX-TS can be applied in a diverse range of application scenarios: In the agriculture domain, it can be used to explain soil moisture forecasts [deforce2024time], where farmers are aiming to optimize their irrigation system. With our method, the system can be tuned to, for example, increase the minimum moisture level. In the context of a smart building digital twin [kreuzer2024ai], the method can provide detailed explanations for CO2 concentration forecasts to, for example, a facility manager, highlighting how different rooms in the building contribute to the overall CO2 levels. The work on hospital occupancy forecasts by [avinash2025time] could be extended with PAX-TS to make use of state-of-the-art forecasters while providing post-hoc explanations for managers.
VI Conclusion
The explanations provided by PAX-TS are both faithful to the forecasting model’s outputs and suitable to address practical questions on time series forecasts from end-users. In comparison with ShapTime and TS-MULE, two state-of-the-art XAI methods, PAX-TS provides both higher and lower granularity explanations. Further, the method is applicable to multivariate data, where it can highlight cross-channel correlations as well as temporal dependency patterns. In our experiments, we classified the identified temporal dependency patterns into six distinct classes. Across all datasets and algorithms, we found that diagonal and bipolar regional dependency patterns tend to perform best. We believe that future work can apply PAX-TS to provide explainability on other datasets and algorithms, which will contribute to our understanding of temporal dependency patterns.
Appendix A Experimental Details
All experiments in the paper were conducted on a single Nvidia A100 GPU, and results were averaged over three random states. All inputs are scaled to have zero mean and unit variance. Table III shows an overview of the forecasting algorithms used in this study, which were used as a basis for PAX-TS as a post-hoc explainability technique.
Algorithm Ref. Approach
Naïve - Naïve
DLinear [zeng2023transformers] MLP
MultiPatchFormer [naghashi2025multiscale] Transformer
SegRNN [lin2023segrnn] RNN
TimeMixer [wang2024timemixer] MLP
iTransformer [liu2023itransformer] Transformer
TimesFM [das2024decoder] LLM
TABLE III: Algorithms evaluated in this study.
Table IV lists the time series datasets used for the experiments in Sec. IV as well as their domain, frequency, and the number of variates. In total, 4 multivariate and 6 univariate datasets are used in the experiments.
Dataset Ref. Domain Frequency Variates
M4 Hourly [makridakis2020m4] Mixed Hourly 1
Weather (a){(a)} Weather Daily 1
Transactions [transactions] Sales Daily 1
CIF [vstvepnivcka2015computational] Finance Monthly 1
Rain (b){(b)} Weather Monthly 1
M4 Yearly [makridakis2020m4] Mixed Yearly 1
Covid (c){(c)} Healthcare Weekly 6
ETTh1 [zhou2021informer] Electricity Hourly 7
Exchange Rate [lai2018modeling] Finance Daily 7
Illness (d){(d)} Healthcare Weekly 7
TABLE IV: Datasets used in this study. (a) https://www.bgc-jena.mpg.de/wetter/, (b) https://www.kaggle.com/datasets/macaronimutton/mumbai-rainfall-data, (c) https://www.kaggle.com/datasets/mexwell/sars-cov-2-germany, (d) https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
Appendix B Local sensitivity analysis and fidelity
We evaluate the local sensitivity of PAX-TS on the CIF dataset with DLinear as the underlying forecasting method by comparing the algorithm’s averaged change ratios, $\bar{r}_{\pi}$, to the model’s own differential behavior. Concretely, we compute Spearman’s $\rho$, cosine similarity, and Kendall’s $\tau$ between $\bar{r}_{\pi}$ and the output Jacobian of DLinear for each horizon timestep of the forecast:
Setup Spearman $\rho$ Cosine similarity Kendall $\tau$
Random -4.19e-4 -8.90e-4 -3.23e-4
PAX-TS 0.778 0.872 0.575
PAX-TS w/o 0.986 0.993 0.926
Gaussian smoothing
TABLE V: Local sensitivity of PAX-TS, compared to random results. We measure Spearman’s rank correlation, cosine similarity, and Kendall’s tau between the change ratios produced by PAX-TS and the Jacobian of the output of DLinear on the CIF dataset, averaged across all samples.
Let a forecasting model be $M_{\theta}:\mathbb{R}^{b}\to\mathbb{R}^{h}$, mapping an input window $\mathbf{x}\in\mathbb{R}^{b}$ to an $h$-step prediction $\hat{\mathbf{y}}=M_{\theta}(\mathbf{x})\in\mathbb{R}^{h}$. PAX-TS returns a timestep-horizon importance matrix $\bar{R}\in\mathbb{R}^{b\times h}$ with entries $\bar{r}_{\pi}$ representing averaged change ratios, as described in Sec. III. We quantify the model’s own differential dependence via the (per-sample) Jacobian
$J[i,j]\;=\;\frac{\partial\hat{y}_{j}}{\partial x_{i}}\quad\text{for}\quad i=1,\dots,b,\;j=1,\dots,h,$
computed by automatic differentiation. For each horizon $j$, define $\mathbf{r}_{j}=\big(\bar{r}_{\pi}[1,j],\dots,\bar{r}_{\pi}[b,j]\big)^{\top}$ and $\mathbf{j}_{j}=\big(J[1,j],\dots,J[b,j]\big)^{\top}$. Based on this, we report rank- and angle-based agreement between $\mathbf{r}_{j}$ and $\mathbf{j}_{j}$:
$\text{Spearman’s }\rho_{j}$ $=\rho\big(\mathrm{rank}(\mathbf{r}_{j}),\,\mathrm{rank}(\mathbf{j}_{j})\big),$
$\text{Cosine}_{j}$ $=\frac{\mathbf{r}_{j}^{\top}\mathbf{j}_{j}}{\|\mathbf{r}_{j}\|_{2}\,\|\mathbf{j}_{j}\|_{2}},$
$\text{Kendall’s }\tau_{j}$ $=\tau\big(\mathbf{r}_{j},\,\mathbf{j}_{j}\big).$
We aggregate across horizons and across a dataset $\{\mathbf{x}^{(n)}\}_{n=1}^{N}$ by simple averaging:
$\overline{\rho}$ $=\frac{1}{Nh}\sum_{n=1}^{N}\sum_{j=1}^{h}\rho_{j}^{(n)},$
$\overline{\text{Cos}}$ $=\frac{1}{Nh}\sum_{n=1}^{N}\sum_{j=1}^{h}\text{Cosine}_{j}^{(n)},$
$\overline{\tau}$ $=\frac{1}{Nh}\sum_{n=1}^{N}\sum_{j=1}^{h}\tau_{j}^{(n)}.$
The resulting coefficients compare PAX-TS’ perturbation-based importance $\bar{r}_{\pi}$ directly to the model’s local derivatives with respect to each input timestep. High $\rho$, cosine, and $\tau$ indicate that the timesteps our method marks as influential are exactly those to which the model is most sensitive. This is the appropriate notion of fidelity here, since PAX-TS explains how the model behaves, not which timesteps are optimal for forecast accuracy (as in LIME/SHAP). When $p(\mathbf{x},\alpha,\phi)$ spreads the perturbation to neighboring timesteps via Gaussian smoothing, $\bar{r}_{\pi}$ incorporates additional context noise. Due to this, rank agreement with the sharp, per-timestep Jacobian may drop slightly. Without smoothing, the perturbation is more localized, yielding the near-perfect alignment reported in Table V. With smoothing (the default in Algorithm 1), correlations remain very high but are modestly affected.
Appendix C Applying PAX-TS
@@
Algorithm 1 shows a step-by-step guide for applying PAX-TS to a forecasting sample, resulting in the change ratios used as explanations. The procedure is the same as described in Sec. III-D. In this section, we additionally highlight that PAX-TS is deterministic, based on the steps in algorithm 1. Given a deterministic forecasting model such as a trained deep learning model which does not introduce any inherent randomness, making a forecast with $M_{\theta}(\cdot)$ is assumed to be deterministic. Further, all permutation methods $p(\cdot)$ introduced in Sec. III are inherently deterministic, as seen from the provided equations. This leads to deterministic change ratios $r_{\pi\alpha}$ and mean change ratios $\bar{r}_{\pi}$, making the outputs of our algorithm deterministic. This is a property that elevates the stability of our method, in comparison to other non-deterministic methods like TS-MULE.
@@
Algorithm 2 details the multivariate cross-channel correlation analysis in a stepwise algorithmic procedure, supporting the reproducibility of this work. With this approach, both multivariate cross-channel correlation heatmaps (as seen in Fig. 7) and cross-channel correlation graphs can be produced (see Fig. 4).
Fig. 6 shows an example of a correlation matrix between different summary statistics of the transactions dataset using iTransformer, as generated by PAX-TS. With this type of explanation, PAX-TS can answer end-user questions on summary statistics, extrema, and trend of the input and forecast. For example, the figure shows a negative correlation between the input minimum and the output’s maximum, mean, and variance. If a user aims to increase the forecast maximum, adjusting the input’s mean, variance, and trend appears to be more effective than increasing its maximum value.
Appendix D Multivariate cross-channel correlations